According to Kepler’s Laws, planets have elliptical orbits, with the sun

at one of the foci. The farthest Pluto gets from the sun is 7.4 billon

kilometers. The closest it gets to the sun


1
Expert's answer
2021-10-26T08:55:02-0400

Let us center the ellipse at "(0,0)" with the major axis along the "x"- axis.

Using the perihelion, "a - c = 4.4" billion km.

Using the aphelion, "a + c = 7.4" billion km.


The semimajor axis is "a = (7.4 + 4.4)\/2 = 5.9" billion km.

"c=7.4-5.9=1.5" billion km.

"b^2=a^2-c^2=(5.9)^2-(1.5)^2=34.81-2.25"

"=32.56" (billion km)2.


The equation of Pluto's orbit assuming a center at "(0,0)" is


"\dfrac{x^2}{a^2}+\dfrac{y^2}{b^2}=1"

"x, y, a," and "b" are in billion km

"\dfrac{x^2}{34.81}+\dfrac{y^2}{32.56}=1"